### v1

# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

Abstract.
For an exact twist map $f$, we introduce an inherent Lipschitz dynamics $\Sigma_{+}$ given by the discrete forward Lax-Oleinik semigroup. We investigate several properties of $\Sigma_{+}$ and show that for any discrete weak K.A.M. solution $u$, the non-differentiable points of $u$ are globally propagated and forward invariant by $\Sigma_{+}$. In particular, such propagating dynamics possesses the same rotation number as the associated Aubry-Mather set with respect to $u$.
A detailed exposition of the corresponding Arnaud’s observation [Arn11] is then provided via $\Sigma_{+}$. Furthermore, we construct and analyze the dynamics on the pseudo-graphs of discrete weak K.A.M. solutions.
Key words and phrases:
twist maps, discrete weak K.A.M. theory, propagation of singularities
2010 Mathematics Subject Classification:
37E10, 37E40, 37J10, 37J30, 37J35
1. Introduction
The celebrated Aubry-Mather theory developed independently by S. Aubry [Aub83, ALD83] and J. Mather [Mat82] provides a novel approach to the study of the dynamics of twist diffeomorphisms of the annulus. The existence of action minimizing sets (referred as Aubry-Mather sets) –which could be invariant circles or invariant Cantor sets etc–is established. Moreover, one may have
  • every Aubry-Mather set is a subset of a homotopically nontrivial circle that is a graph over the zero section,
  • every Aubry-Mather set is associated a rotation number and
  • all the Aubry-Mather sets are vertically ordered in the annulus according to the rotation numbers. See [Gol01].
We concentrate on those Aubry-Mather sets with gaps and intend to study how these gaps evolve. To this aim, we will introduce a singular dynamics such that the gaps are forward invariant by such dynamics.
Let $f$ be an exact twist map on the annulus $\mathbb{T}\times\mathbb{R}$ and $F$ be its lift on its universal cover $\mathbb{R}^{2}$. Because of the exactness of $f$, one can associate $F$ with a $C^{2}$ generating function $S:\mathbb{R}^{2}\rightarrow\mathbb{R}$.
Let $c\in\mathbb{R}$ be some cohomology class [^1] In the present paper, the cohomology class $c$ is fixed at the beginning and one can assume $c=0$ for simplicity. The reason why we insist writing $c$ explicitly in the introduction is that we will apply the results here for future work. of $\mathbb{T}$. One can also define the lift $F_{c}$ and the corresponding generating function $S_{c}$ at cohomology class $c$ given by
$F_{c}(x,p)=F(x,p+c)-(0,c)\text{ and }S_{c}(x,y)=S(x,y)+c(x-y)\text{ respectively}.$
We then define respectively the discrete backward and forward Lax-Oleinik operators $T^{-}$ and $T^{+}$ by
$T^{-}[u](x):=$ $\inf_{y\in\mathbb{R}}\big{\{}u(y)+S_{c}(y,x)\big{\}},\qquad\forall x\in\mathbb{R}$
$T^{+}[u](x):=$ $\sup_{y\in\mathbb{R}}\big{\{}u(y)-S_{c}(x,y)\big{\}},\qquad\forall x\in\mathbb{R}$
for every continuous $1$-periodic function $u:\mathbb{R}\rightarrow\mathbb{R}$. Then one can prove that there exists a unique constant $\overline{S}_{c}\in\mathbb{R}$ such that the equation $u=T^{-}[u]-\overline{S}_{c}$ admits at least one solution $u_{c}$ (see for example [GT11, ST18]). This min-plus eigenvalue problem in ergodic optimization gives rise to the following two notions in celebrated Mather-Fathi theory (see [Mat91, Fat97]) for Tonelli Lagrangian dynamics:
  • Mather’s $\alpha$-function: $H^{1}(\mathbb{T},\mathbb{R})\rightarrow\mathbb{R},c\mapsto-\bar{S}_{c}$, which is differentiable everywhere.
  • a discrete weak K.A.M. solution $u_{c}$ for $f$ at cohomology class $c$, which is semi-concave and so differentiable almost everywhere.
Therefore, we could state our first result below:
Theorem 1.1.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the map $\Sigma_{+}:\mathbb{R}\rightarrow\mathbb{R},\ x\mapsto y$ is well-defined and implicitly given by
$T^{+}[u_{c}](x)=u_{c}(y)-S_{c}(x,y).$
Moreover, the map $\Sigma_{+}$ satisfies the following properties:
  1. (i)
    $\Sigma_{+}$ is a lift of a circle map with degree $1$, i.e., $\Sigma_{+}(x+1)=\Sigma_{+}(x)+1$;
  2. (ii)
    $\Sigma_{+}$ is non-decreasing and Lipschitz continuous;
  3. (iii)
    $\Sigma_{+}$ propagates singularities of $u_{c}$, i.e., if x is not a differentiable point of $u_{c}$, neither is $\Sigma_{+}(x)$; for simplicity, we denote by $\operatorname{Sing}(u_{c})$ the set of all the non-differentiable points of $u_{c}$;
  4. (iv)
    $\Sigma_{+}$ admits a rotation number for positive iterations, that is,
    $\rho(\Sigma_{+}):=\lim_{n\to+\infty}\frac{1}{n}\left(\Sigma_{+}^{n}(x)-x\right)$
    exists for all $x\in\mathbb{R}$ and is independent of $x$. In fact, $\rho(\Sigma_{+})=\alpha^{\prime}(c)$.
As an application of the above properties of the singular dynamics, we immediately obtain that the singularities of a discrete weak K.A.M. solution $u$ can be globally propagated with a rotation number.
Theorem 1.2.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Let $x_{0}\in\operatorname{Sing}(u_{c})$ and one can define the one-sided sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ iteratively by:
$x_{n+1}:=\mathop{\arg\max}_{y\in\mathbb{R}}\{u_{c}(y)-S_{c}(x_{n},y)\}\qquad\text{ for any }n\in\mathbb{Z}_{+}.$
Then, we have
  • (a)
    the whole one-sided sequence $\{x_{n}\}_{n\in\mathbb{N}}\subset\operatorname{Sing}(u_{c})$;
  • (b)
[…]
  • (c)
    $\rho_{c}$ equals to the rotation number $\alpha^{\prime}(c)$ of the associated Aubry-Mather set.
Note that item (a) of Theorem 1.2 could be regarded as a discrete version of [CC17, Lemma 3.2] (see also [Zha20]). To the best of our knowledge, items (b),(c) are new in the literature of twist maps.
As an immediate application, we now investigate the relation between the number of the singular set $\operatorname{Sing}(u_{c})$ of a given discrete weak K.A.M. solution $u_{c}$ and the arithmetic property of the associated rotation number.
If $\operatorname{Sing}(u_{c})\neq\emptyset$, we can pick $x_{0}\in\operatorname{Sing}(u_{c})$ and define $x_{n+1}:=\Sigma_{+}(x_{n}),\ \forall n\in\mathbb{Z}_{+}$. In addition, if $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is a finite set, we know that there exist two different natural numbers $n<m$ such that $x_{m}-x_{n}\in\mathbb{Z}$. Then we obtain that $\rho_{c}=(x_{m}-x_{n})/(m-n)\in\mathbb{Q}$. Hence $\alpha^{\prime}(c)$ also belongs to $\mathbb{Q}$.
On the other hand, if $\alpha^{\prime}(c)\in\mathbb{R}\setminus\mathbb{Q}$, due to the fact $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is at most countable (see [CS04, Proposition 4.1.3]), we conclude
Corollary 1.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$.
  • (1)
    If $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is not empty and has finite elements, then $\alpha^{\prime}(c)\in\mathbb{Q}$.
  • (2)
    If $\alpha^{\prime}(c)\in\mathbb{R}\setminus\mathbb{Q}$, then $\operatorname{Sing}(u_{c})/\mathbb{Z}$ is either empty or infinitely countable.
In the next, we will use the singular dynamics $\Sigma_{+}$ to give a detailed explanation of Arnaud’s observation for the connection between pseudo-graph and the Lax-Oleinik semi-group [Arn11] in the case of twist maps (see also [Zav23]).
Theorem 1.3.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the graphs $\operatorname{Graph}(\Sigma_{+}),\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$ and $\operatorname{Graph}(c+\nabla^{+}u_{c})$ are Lipschitz submanifolds of $\mathbb{R}^{2}$ associated by the commutative diagram:
{tikzcd}
where $\Gamma^{+}(x,y)=(y,\partial_{2}S(x,y))$, $\Gamma^{-}(x,y)=(x,-\partial_{1}S(x,y))$ and $F$ is the lift of $f$ associated to $S$.
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
Figure 1. The graphs $\operatorname{Graph}(\Sigma_{+}),\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$ and $\operatorname{Graph}(c+\nabla^{+}u_{c})$ at cohomology class $c=0$ in the pendulum case.
Following Arnaud and Zavidovique’s article [AZ23], we call $\operatorname{Graph}(c+\nabla^{+}u_{c})$ the full pseudo-graph of a weak K.A.M. solution $u_{c}$ at cohomology class $c$. If $x\in\operatorname{Sing}(u_{c})$, then $\nabla^{+}u_{c}(x)$ is not a singleton and we refer to $\{x\}\times(c+\nabla^{+}u_{c}(x))$ as a vertical segment of the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$. Notice that there is a natural forward map
$\Sigma_{+}\times\Sigma_{+}:\operatorname{Graph}(\Sigma_{+})$ $\rightarrow\operatorname{Graph}(\Sigma_{+})$
$(x,\Sigma_{+}(x))$ $\mapsto(\Sigma_{+}(x),\Sigma_{+}^{2}(x)).$
Hence for any cohomology class $c$, there are two maps
$\Omega_{c}:\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$ $\rightarrow\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}]),$
$\Lambda_{c}:\operatorname{Graph}(c+\nabla^{+}u_{c})$ $\rightarrow\operatorname{Graph}(c+\nabla^{+}u_{c})$
induced by the diagram:
{tikzcd}
The dynamical systems defined by $\Lambda_{c}$ and $\Omega_{c}$ are conjugate and also termed as the singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$ and $\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$ respectively (see Figures 2 and 3 for an illustration).
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
Figure 2. Singular dynamics on the full pseudo-graph $\operatorname{Graph}(c+\nabla^{+}u_{c})$ of a discrete weak K.A.M. solution $u_{c}$ at cohomology class $c$, where different colors are used to represent the corresponding relationships before and after the map $\Lambda_{c}$.
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
Figure 3. Singular dynamics on the $\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u_{c}])$, where different colors are used to represent the corresponding relationships before and after the map $\Omega_{c}$.
Consequently, one could analyze the dynamics $\Lambda_{c}$ on the full pseudo-graphs $\operatorname{Graph}(c+\nabla^{+}u_{c})$ and has the following theorem.
Theorem 1.4.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the map
$\Lambda_{c}:\operatorname{Graph}(c+\nabla^{+}u_{c})$ $\rightarrow\operatorname{Graph}(c+\nabla^{+}u_{c})$
$(x,p)$ $\mapsto(y,\partial_{2}S(x,y))$
is well-defined and implicitly given by
$T^{+}[u_{c}](x)=u_{c}(y)-S_{c}(x,y).$
Moreover, the map $\Lambda_{c}$ satisfies the following properties:
  1. (i)
    $\Lambda_{c}(x+1,p)=\Lambda_{c}(x,p)+(1,0)$;
  2. (ii)
    $\Lambda_{c}$ is Lipschitz continuous;
  3. (iii)
    $\Lambda_{c}$ does not reverse the orientation of the full pseudo-graph.
  4. (iv)
    the image of a vertical segment under $\Lambda_{c}$ is a point that belongs to a vertical segment.
  5. (v)
    $\Lambda_{c}$ admits a rotation number for positive iterations, that is,
    $\rho(\Lambda_{c}):=\lim_{n\to+\infty}\frac{1}{n}\left(\pi_{1}\circ\Lambda_{c}^{n}(x,p)-x\right)$
    exists for all $(x,p)\in\operatorname{Graph}(c+\nabla^{+}u_{c})$ and is independent of $(x,p)$. In fact, $\rho(\Lambda_{c})=\alpha^{\prime}(c)$.
Observing Figure 2 reveals that within the singular dynamics, certain points are attracted to the singularities. However, it is also evident that some points do not propagate into the singularities during the iteration of $\Lambda_{c}$. Thus, one could conclude:
Theorem 1.5.
Let $f$ be a twist map of the annulus and $c\in\mathbb{R}$ be some cohomology class of $\mathbb{T}$. Let $u_{c}:\mathbb{R}\rightarrow\mathbb{R}$ be a discrete weak K.A.M. solution for $f$ at cohomology class $c$. Then the two sets
$c+\mathcal{I}^{*}(u_{c}):=\bigcap_{n\geq 0}F^{-n}(\overline{\operatorname{Graph}(c+\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u_{c})});$
$c+\mathcal{K}^{*}(u_{c}):=\bigcup_{\begin{subarray}{c}x\in\operatorname{Sing}(u_{c})\\ p\in c+\nabla^{+}u_{c}(x)\end{subarray}}\Lambda_{c}(x,p)$
satisfy the following properties
  1. (i)
    the set $c+\mathcal{I}^{*}(u_{c})$ is closed, non-empty and $\Lambda_{c}$-invariant;
  2. (ii)
    the restriction of $\Lambda_{c}$ on $c+\mathcal{I}^{*}(u_{c})$ is the lift $F$ of the twist map $f$;
  3. (iii)
    the set $c+\mathcal{K}^{*}(u_{c})$ is at most countable and contained within the vertical segments of $\operatorname{Graph}(c+\nabla^{+}u_{c})$;
  4. (iv)
    for any $(x,p)\in\operatorname{Graph}(c+\nabla^{+}u_{c})\setminus(c+\mathcal{I}^{*}(u_{c}))$, there exists $n\in\mathbb{Z}_{+}$ such that
    $\Lambda_{c}^{n}(x,p)\in c+\mathcal{K}^{*}(u_{c}).$
Before ending this section, as an application of $c+\mathcal{I}^{*}(u_{c})$, we represent the Aubry set in Theorem 2.15 that:
$c+\mathcal{A}^{*}_{c}=\bigcap\left(c+\mathcal{I}^{*}(u_{c})\right),$
where the intersection is considered over all discrete weak K.A.M. solutions $u_{c}$ at the cohomology class $c$. This constitutes the weak Hamilton-Jacobi theorem (see e.g. [Fat08]) in the discrete setting. For the sake of simplicity, all proofs in Sections 3-5 only deals with the case where $c=0$.
2. Preliminary
Notations.
  • $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ is the $d$-dimensional torus and its cotangent bundle $T^{*}\mathbb{T}^{d}=\mathbb{T}^{d}\times(\mathbb{R}^{d})^{*}$ is the $2d$-dimensional annulus. The points of the annulus are denoted by $(\theta,p)\in\mathbb{T}^{d}\times(\mathbb{R}^{d})^{*}$ and the points of its universal cover are denoted by $(x,p)\in\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}$.
  • Let $S(x_{1},\dots,x_{2d})$ be a $C^{1}$ real-valued function defined on $\mathbb{R}^{2d}$. We denote by $\partial_{1}S$ and $\partial_{2}S$ the $(\mathbb{R}^{d})^{*}$-valued functions $(\partial S/\partial x_{1},\dots,\partial S/\partial x_{d})$ and $(\partial S/\partial x_{d+1},\dots,\partial S/\partial x_{2d})$ respectively.
  • For the product space $X\times Y$, the maps $\pi_{1}:X\times Y\rightarrow X$ and $\pi_{2}:X\times Y\rightarrow Y$ are the projections on the first and second variable.
  • Let $\mathcal{G}:X\rightarrow 2^{Y}$ be a set-valued map, we define its graph as a subset of $X\times Y$:
    $\operatorname{Graph}(\mathcal{G}):=\{(x,y)\in X\times Y\mid y\in\mathcal{G}(x)\}.$
    For example, if $u:\mathbb{T}^{d}\rightarrow\mathbb{R}$ is semiconcave, we will denote by $\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$ the partial graph of $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u$:
    $\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)=\{(x,y)\in\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}\mid u\text{ is differentiable at $x$ and }y=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)\}.$
    And we will denote by $\operatorname{Graph}(\nabla^{+}u)$ the full pseudo-graph of $u$:
    $\operatorname{Graph}(\nabla^{+}u)=\{(x,y)\in\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}\mid y\in\nabla^{+}u(x)\}.$
2.1. Discrete weak K.A.M. theory on $d$-dimensional torus
Let us recall some fundamental definitions and properties on semiconcave functions (see [CS04, Chapter 3] for instance).
Definition 2.1.
Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function.
  • For any $x\in\mathbb{R}^{d}$, the sets
    $\nabla^{-}u(x)=\left\{p\in(\mathbb{R}^{d})^{*}:\liminf_{y\rightarrow x}\frac{u(y)-u(x)-p(y-x)}{\|y-x\|}\geq 0\right\}$
    $\nabla^{+}u(x)=\left\{p\in(\mathbb{R}^{d})^{*}:\limsup_{y\rightarrow x}\frac{u(y)-u(x)-p(y-x)}{\|y-x\|}\leq 0\right\}$
    are called, respectively, the superdifferential and subdifferential of $u$ at $x$. From the definition it follows that, for any $x\in\mathbb{R}^{d}$, $-\nabla^{-}u(x)=\nabla^{+}(-u)(x).$
  • Let $\Omega\subset\mathbb{R}^{d}$ be a convex open set, we say that $u$ is semiconcave (with a linear modulus) on $\Omega$ if there exists a constant $C\geq 0$ such that
[…]
    for any $x,y\in\Omega$ and $\lambda\in[0,1]$. Then $C$ is called the semiconcavity constant of $u$. A function $v$ is called semiconvex on $\Omega$ if $-v$ is semiconcave. We say that $u$ is locally semiconcave (with a linear modulus) if for each $x\in\mathbb{R}^{d}$, there exists a convex open neighborhood $U_{x}$ such that $u$ is semiconcave on $U_{x}$.
  • For any $x\in\mathbb{R}^{d}$, a covector $p\in(\mathbb{R}^{d})^{*}$ is called a reachable differential of $u$ at $x$ if a sequence $\{x_{n}\}\subset\mathbb{R}^{d}\backslash\{x\}$ exists such that $u$ is differentiable at $x_{k}$ for each $k\in\mathbb{Z}_{+}$, and
    $\lim_{k\rightarrow\infty}x_{k}=x,\quad\lim_{k\rightarrow\infty}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u\left(x_{k}\right)=p.$
    The set of all reachable gradients of $u$ at $x$ is denoted by *$\nabla^{*}u(x)$.
Proposition 2.2.
  1. (i)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be semiconcave function on $\Omega$ with linear modulus and constant $C$. Then
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}|x-x_{0}|^{2}$
    for any $x,x_{0}\in\Omega$ and $p_{0}\in\nabla^{+}u(x_{0})$.
  2. (ii)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function, $\Omega\subset\mathbb{R}^{d}$ be a convex open set and $x_{0}\in\Omega$. If there exists $C\geq 0$ and $p_{0}\in(\mathbb{R}^{d})^{*}$ such that
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}|x-x_{0}|^{2},\ \quad\forall x\in\Omega$
    then $p_{0}\in\nabla^{+}u(x_{0})$.
  3. (iii)
    Let $u,v:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be two continuous functions, $\Omega\subset\mathbb{R}^{d}$ be a convex open set and $x\in\Omega$. If $u\leq v$ on $\Omega$ and $u(x)=v(x)$, then $\nabla^{+}u(x)\supset\nabla^{+}v(x)$ and $\nabla^{-}u(x)\subset\nabla^{-}v(x)$.
  4. (iv)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function and $x\in\mathbb{R}^{d}$. Then $u$ is differentiable at $x$ if and only if $\nabla^{+}u(x)$ and $\nabla^{-}u(x)$ are both nonempty; in this case we have that
    $\nabla^{+}u(x)=\nabla^{-}u(x)=\{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)\}.$
  5. (v)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be semiconcave function on $\Omega$. Then *$\nabla^{*}u(x)\subset\partial\nabla^{+}u(x)$ for all $x\in\Omega$, where $\partial$ is the boundary taken with respect to the standard topology of $\mathbb{R}^{d}$.
Hypothesis 1.
The function $S:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the following properties:
  1. (1)
    $S$ is of class $C^{1}$;
  2. (2)
    $S$ is coercive:
    $\lim_{\|x-y\|\rightarrow+\infty}S(x,y)=+\infty;$
  3. (3)
    $S$ is locally semiconcave;
  4. (4)
    $S$ is diagonal-periodic:
    $\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},m\in\mathbb{Z}^{d}\quad S(x+m,y+m)=S(x,y);$
  5. (5)
    $S$ is ferromagnetic; that is, the two maps $x\mapsto\partial_{2}S(x,y)$ and $y\mapsto\partial_{1}S(x,y)$ are homeomorphisms between $\mathbb{R}^{d}$ and $(\mathbb{R}^{d})^{*}$ for all $(x,y)$.
Remark 2.3.
The notion of ferromagnetic terminology is introduced by Garibaldi and Thieullen (see [GT11, Definition 2.4]). If $S$ satisfies Hypothesis 1, then the discrete standard map $F:\mathbb{R}^{d}\times(\mathbb{R}^{d})^{*}\righttoleftarrow$ can be defined by
$F(x,p)=(y,p^{\prime}):\Longleftrightarrow\begin{cases}p=-\partial_{1}S(x,y)\\ p^{\prime}=\partial_{2}S(x,y)\end{cases}.$
The map $F$ is a homeomorphism and can be seen as a discrete version of the Hamiltonian flow.
We recall the discrete backward Lax-Oleinik semigroup and discrete forward Lax-Oleinik semigroup defined in the introduction:
$\forall x\in\mathbb{R}^{d},\quad T^{-}[u](x):=$ $\inf_{y\in\mathbb{R}^{d}}\big{\{}u(y)+S(y,x)\big{\}},$
$\forall x\in\mathbb{R}^{d},\quad T^{+}[u](x):=$ $\sup_{y\in\mathbb{R}^{d}}\big{\{}u(y)-S(x,y)\big{\}},$
for every continuous $\mathbb{Z}^{d}$-periodic function $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (namely $u\in C^{0}(\mathbb{T}^{d})$).
Remark 2.4.
  1. (a)
    The image $T^{-}(C^{0}(\mathbb{T}^{d}))$ of $T^{-}$ is equicontinuous, as shown in [GT11, Lemma 4.2]. Moreover, it constitutes a family of semiconcave functions with linear modulus and the same constant, as demonstrated in [Zav23, Proposition 2.4.10].
  2. (b)
    [Zav23, Proposition 1.3.6] For any $u\in C^{0}(\mathbb{T}^{d})$ and $n\in\mathbb{Z}_{+}$,
    $T^{+n}\circ T^{-n}[u]\leq u\leq T^{-n}\circ T^{+n}[u].$
  3. (c)
    As $S$ is continuous and coercive, for any $x\in\mathbb{R}^{d}$, we can find $y\in\mathbb{R}^{d}$ that attains the infimum in the definition of $T^{-}[u](x)$ and the supremum in the definition of $T^{+}[u](x)$. In other words, given $u\in C^{0}(\mathbb{T}^{d})$, for each $y\in\mathbb{R}^{d}$, there exists $x\in\mathbb{R}^{d}$ such that
    $T^{-}[u](y)=u(x)+S(x,y)$
    and for each $x\in\mathbb{R}^{d}$, there exists $y\in\mathbb{R}^{d}$ such that
    $T^{+}[u](x)=u(y)-S(x,y).$
In the next, we will characterize the regularity of $u,T^{-}[u],T^{+}[u]$ in the following Lemmas 2.5, 2.6 and 2.12 whose proofs are given in the appendix.
Lemma 2.5.
Assume that $S$ satisfies Hypothesis 1. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
    If $T^{-}[u](y)=u(x)+S(x,y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}u(x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}T^{-}[u](y);$
  2. (b)
    if $T^{+}[u](x)=u(y)-S(x,y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}T^{+}[u](x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}u(y);$
Lemma 2.6.
Assume that $S$ satisfies Hypothesis 1. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
    Let $y\in\mathbb{R}^{d}$. Then there exists a unique $x\in\mathbb{R}^{d}$ such that $T^{-}[u](y)=u(x)+S(x,y)$ if and only if $T^{-}[u]$ is differentiable at $y$.
  2. (b)
    Let $x\in\mathbb{R}^{d}$. Then there exists a unique $y\in\mathbb{R}^{d}$ such that $T^{+}[u](x)=u(y)-S(x,y)$ if and only if $T^{+}[u]$ is differentiable at $x$.
Theorem 2.7 (Discrete Weak K.A.M.).
There exists a unique constant $\overline{S}$ such that the equation $u=T^{-}[u]-\overline{S}$ admits at least one solution $u\in C^{0}(\mathbb{T}^{d})$.
Such functions $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are called discrete weak K.A.M solutions. The constant $\bar{S}$ is called the effective interaction.
Definition 2.8.
A function $u\in C^{0}(\mathbb{T}^{d})$ is called a sub-action with respect to $S$ if
$\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},\ \quad u(y)-u(x)\leq S(x,y)-\bar{S},$
or equivalently, if $u\leq T^{-}[u]-\bar{S}$, or equivalently, if $u\geq T^{+}[u]+\bar{S}$.
Remark 2.9.
For any sub-action $u$, the functions $T^{\mp}[u]$ are also sub-actions [Zav23, Proposition 1.2.6]. Thus, we have
$u\leq T^{-}[u]-\bar{S}\leq T^{-2}[u]-2\bar{S}\leq T^{-3}[u]-3\bar{S}\leq\cdots$
This implies that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is non-decreasing. Moreover, we can show that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is uniformly bounded [Zav23, Proposition 1.2.3]. Note that (a) of Remark 2.4 states that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is equicontinuous. Therefore, $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $u_{-}\in C^{0}(\mathbb{T}^{d})$, which is a discrete weak K.A.M. solution.
Definition 2.10.
Let $u$ be a sub-action. We say that $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is $u$-calibrated if $u(y)-u(x)=S(x,y)-\bar{S}$. A sequence $\{x_{n}\}_{n\in\mathbb{Z}}$ is called $u$-calibrated if $(x_{n},x_{n+1})$ is $u$-calibrated for all $n\in\mathbb{Z}$. The non-strict set $\mathcal{NS}(u)$ of $u$ collects all $u$-calibrated pairs:
$\mathcal{NS}(u):=\{(x,y)\in\mathbb{R}^{2}\mid u(y)-u(x)=S(x,y)-\bar{S}\}.$
Remark 2.11.
Let $u$ be a sub-action. If $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is $u$-calibrated, then we have
  • $u(y)=T^{-}[u](y)-\bar{S}$;
  • $u(x)=T^{+}[u](x)+\bar{S}$;
  • $T^{-}[u](y)=u(x)+S(x,y)$;
  • $T^{+}[u](x)=u(y)-S(x,y)$.
Lemma 2.12.
Assume that $S$ satisfies Hypothesis 1. Let $u$ be a sub-action.
  1. (a)
    If $(x,y)\in\mathcal{NS}(T^{-}[u])$, then both $u$ and $T^{-}[u]$ are differentiable at $x$. Moreover, $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=dT^{-}[u](x)=-\partial_{1}S(x,y)$.
  2. (b)
    If $(x,y)\in\mathcal{NS}(T^{+}[u])$, then both $u$ and $T^{+}[u]$ are differentiable at $y$. Moreover, $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)=dT^{+}[u](y)=\partial_{2}S(x,y)$.
Definition 2.13.
The projected Aubry set is defined as
$\mathcal{A}:=\pi_{1}\left(\bigcap_{u\text{ is a sub-action}}\mathcal{NS}(u)\right).$
This may not be the most common definition, but it can be proven to be equivalent to other definitions (see [GT11, Proposition 10.5]). We list below, without proof, some properties of the projected Aubry set $\mathcal{A}$. Interested readers may refer to Garibaldi and Thieullen’s paper [GT11] and Zavidovique’s note [Zav23].
Proposition 2.14.
Assume that $S$ satisfies Hypothesis 1.
  1. (i)
    $\mathcal{A}$ is closed, non-empty and
    $\mathcal{A}=\pi_{2}\left(\bigcap_{u\text{ is a sub-action}}\mathcal{NS}(u)\right);$
  2. (ii)
    $\mathcal{A}+m=\mathcal{A},\ \forall m\in\mathbb{Z}^{d}$;
  3. (iii)
    Any sub-action $u$ is differentiable on $\mathcal{A}$. Moreover, the differential $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u:\mathcal{A}\rightarrow(\mathbb{R}^{d})^{*}$ is independent of $u$;
  4. (iv)
    For any $x_{0}\in\mathcal{A}$, there exists a unique sequence $\{x_{n}\}_{n\in\mathbb{Z}}$ which is calibrated by any sub-action $u$;
  5. (v)
    Denote the Aubry set $\mathcal{A}^{*}=\{(x,p)\mid p=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x),x\in\mathcal{A}\}$. Then the projection $\pi_{1}:\mathcal{A}^{*}\rightarrow\mathcal{A}$ is a bi-Lipschitz homeomorphism.
  6. (vi)
    $\mathcal{A}^{*}$ is an F-invariant set, i.e., $F(\mathcal{A}^{*})=\mathcal{A}^{*}$ and $F^{-1}(\mathcal{A}^{*})=\mathcal{A}^{*}$.
Theorem 2.15.
Assume that $S$ satisfies Hypothesis 1. If $u$ is a discrete weak K.A.M. solution, then we have
  • (1)
    For any $n\geq 1$,
    $F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
    So one can define the closed, non-empty and $F$-invariant set
    *$\mathcal{I}^{*}(u):=\bigcap_{n\geq 0}F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right).$
  • (2)
    there exists a discrete weak K.A.M. solution $\widetilde{u}_{-}$ such that $\mathcal{A}^{*}=\mathcal{I}^{*}(\widetilde{u}_{-}).$
  • (3)
    moreover,
    $\mathcal{A}^{*}=\bigcap\mathcal{I}^{*}(u),$
    where the intersection is taken over all discrete weak K.A.M. solutions.
Proof.
(1). To show that $F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$ for all $n\geq 1,$ it suffices to demonstrate
$F^{-1}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
Consider a point $(y,p^{\prime})\in\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}$. There exists a sequence $\{(y_{k},p^{\prime}_{k})\}_{k\in\mathbb{Z}_{+}}\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$ that converges to $(y,p^{\prime})$ as $k\rightarrow+\infty$. If $(y_{k},p^{\prime}_{k})\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$, then $u$ is differentiable at $y_{k}$ and $p^{\prime}_{k}=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y_{k})$. Since $u=T^{-}[u]-\bar{S}$, the function $T^{-}[u]$ is also differentiable at $y_{k}$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{-}[u](y_{k})=p^{\prime}_{k}$. Using item (a) of Lemma 2.6, there exists a unique $x_{k}\in\mathbb{R}^{d}$ such that $T^{-}[u](y_{k})=u(x_{k})+S(x_{k},y_{k})$. By item (a) of Lemma 2.5, we have
$\partial_{2}S(x_{k},y_{k})=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{-}[u](y_{k})=p^{\prime}_{k}.$
Since $S$ is is ferromagnetic, the convergence $(y_{k},p_{k}^{\prime})\rightarrow(y,p^{\prime})$ implies that $\{x_{k}\}_{k\in\mathbb{Z}_{+}}$ converges to some $x\in\mathbb{R}^{d}$ and
$\partial_{2}S(x,y)=p^{\prime}.$
Then, from $T^{-}[u](y_{k})=u(x_{k})+S(x_{k},y_{k})$, we get
$T^{-}[u](y)=u(x)+S(x,y).$
Using $u=T^{-}[u]-\bar{S}$ again, we have
$u(y)+\bar{S}=T^{-}[u](y)=u(x)+S(x,y).$
That is, $(x,y)\in\mathcal{NS}(u)=\mathcal{NS}(T^{-}[u])$. Then it follows from the item (a) of Lemma 2.12 that $u$ is differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=-\partial_{1}S(x,y)$. Combine $\partial_{2}S(x,y)=p^{\prime}$ and $-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)$, by Remark 2.3, we have
$F^{-1}(y,p^{\prime})=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)).$
Therefore, $F^{-1}(y,p^{\prime})\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$.
(2). Let $\widetilde{u}$ be a $C^{1,1}$ sub-action such that $\mathcal{A}=\pi_{1}(\mathcal{NS}(\widetilde{u}))=\pi_{2}(\mathcal{NS}(\widetilde{u}))$ in [Zav10]. Thanks to Remark 2.9, we can define
$\widetilde{u}_{-}:=\lim_{n\to+\infty}T^{-n}[\widetilde{u}]-n\bar{S}.$
Consider a point $(x_{0},p_{0})\in\mathcal{I}^{*}(\widetilde{u}_{-})$ and denote $(x_{k},p_{k}):=F^{k}(x_{0},p_{0}),\ \forall k\in\mathbb{Z}_{+}$. Then we know that $\widetilde{u}_{-}$ is differentiable at $x_{k}$ with differential $p_{k}$ for each $k\geq 1$. Using item (a) of Lemma 2.6, there exists a unique $w_{k}\in\mathbb{R}^{d}$ such that $T^{-}[\widetilde{u}_{-}](x_{k})=\widetilde{u}_{-}(w_{k})+S(w_{k},x_{k})$ for each $k\geq 1$. By item (a) of Lemma 2.5, we have
$\partial_{2}S(w_{k},x_{k})=p_{k},\quad\forall k\geq 1.$
Since $S$ is ferromagnetic and $F(x_{k-1},p_{k-1})=(x_{k},p_{k})$, we have $x_{k-1}=w_{k}$ for each $k\geq 1$. In particular,
$\widetilde{u}_{-}(x_{k})+\bar{S}=T^{-}[\widetilde{u}_{-}](x_{k})=\widetilde{u}_{-}(x_{k-1})+S(x_{k-1},x_{k}),\quad\forall k\geq 1.$
That is, the pair $(x_{k-1},x_{k})$ is $\widetilde{u}_{-}$-calibrated for each $k\geq 1$. Then, by Remark 2.11,
$T^{+}[\widetilde{u}_{-}](x_{k})-T^{+}[\widetilde{u}_{-}](x_{k-1})$
$=$ $\left(\widetilde{u}_{-}(x_{k})-\bar{S}\right)-\left(\widetilde{u}_{-}(x_{k})-S(x_{k-1},x_{k})\right)$
$=$ $S(x_{k-1},x_{k})-\bar{S}.$
That is, the pair $(x_{k-1},x_{k})$ is $T^{+}[\widetilde{u}_{-}]$-calibrated for each $k\geq 1$. We claim, by induction, that for all $k,n\in\mathbb{Z}_{+}$, the pair $(x_{k},x_{k+1})$ is $T^{+n}[\widetilde{u}_{-}]$-calibrated. Specifically, $(x_{0},x_{1})$ is $T^{+n}[\widetilde{u}_{-}]$-calibrated for each $n\in\mathbb{Z}_{+}$. Utilizing Remark 2.11 once again, we obtain
$T^{+n}[\widetilde{u}_{-}](x_{0})+n\bar{S}=\widetilde{u}_{-}(x_{0}),\quad\forall n\in\mathbb{Z}_{+}.$
Let $\varepsilon(n):=\sup\left(\widetilde{u}_{-}-T^{-n}[\widetilde{u}]+n\bar{S}\right)$. Then, based on item (b) of Remark 2.4, for each $n\in\mathbb{Z}_{+}$, we have
$\widetilde{u}_{-}-\widetilde{u}$ $\leq\widetilde{u}_{-}-T^{+n}\circ T^{-n}[\widetilde{u}]$
$\leq\widetilde{u}_{-}-T^{+n}[\widetilde{u}_{-}+n\bar{S}-\varepsilon(n)]$
$=\left(\widetilde{u}_{-}-T^{+n}[\widetilde{u}_{-}]-n\bar{S}\right)+\varepsilon(n).$
Setting $x=x_{0}$ and letting $n\to+\infty$, we obtain $\widetilde{u}_{-}(x_{0})-\widetilde{u}(x_{0})\leq 0$. Given that $\widetilde{u}\leq\widetilde{u}_{-}$, we conclude $\widetilde{u}(x_{0})=\widetilde{u}_{-}(x_{0})$. Recall from Remark 2.9 that
$\widetilde{u}\leq T^{-}[\widetilde{u}]-\bar{S}\leq\cdots\leq\widetilde{u}_{-}.$
Thus, $T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{0})+\bar{S}$. Applying (c) of Remark 2.4, we find $x_{-1}\in\mathbb{R}^{d}$ such that
$T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{-1})+S(x_{-1},x_{0})$
This leads to
$\widetilde{u}(x_{0})+\bar{S}=T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{-1})+S(x_{-1},x_{0}).$
Therefore, $(x_{-1},x_{0})\in\mathcal{NS}(\widetilde{u})$, implying $x_{0}\in\pi_{2}(\mathcal{NS}(\widetilde{u}))=\mathcal{A}$. With $p_{0}=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\widetilde{u}_{-}(x_{0})$, we have $(x_{0},p_{0})\in\mathcal{A}^{*}$. Consequently, *$\mathcal{I}^{*}(\widetilde{u}_{-})\subset\mathcal{A}^{*}.$ We have proved that $\mathcal{A}^{*}\subset\mathcal{I}^{*}(u)$ for each discrete weak K.A.M. solution $u$. Therefore, $\mathcal{A}^{*}=\mathcal{I}^{*}(\widetilde{u}_{-}).$
(3). It follows from items (iii) and (vi) of Proposition 2.14 that $\mathcal{A}^{*}\subset\mathcal{I}^{*}(u)$ for each discrete weak K.A.M. solution $u$. Thus, the desired result can be immediately deduced from item (2). ∎
2.2. Twist maps of the $2$-dimensional annulus
We begin by introducing the notion of twist maps.
Definition 2.16.
A twist map of the $2$-dimensional annulus is a $C^{1}$-diffeomorphism $f:\mathbb{T}\times\mathbb{R}^{*}\righttoleftarrow$ such that
  1. (1)
    $f$ is isotopic to identity ;
  2. (2)
    $f$ is exact symplectic, i.e. if $f(\theta,p)=(\theta^{\prime},p^{\prime})$, then the $1$-form $p^{\prime}d\theta^{\prime}-pd\theta$ is exact;
  3. (3)
    $f$ twists verticals to the right: if $F:\mathbb{R}\times\mathbb{R}^{*}\righttoleftarrow$, denoted by $F(x,p)=(y,p^{\prime})$, is a lift of $f$ then for all $x\in\mathbb{R}$, then there exists $\varepsilon>0$ such that
    $0<\frac{\partial y}{\partial p}(x,p)<\frac{1}{\varepsilon},\quad\forall(x,p)\in\mathbb{R}\times\mathbb{R}^{*}.$
Definition 2.17.
A generating function of the lift $F:\mathbb{R}\times\mathbb{R}^{*}\righttoleftarrow$ is a $C^{2}$ function $S:\mathbb{R}^{2}\rightarrow\mathbb{R}$ such that
  1. (1)
    $S$ is diagonal-periodic:
    $\forall(x,y)\in\mathbb{R}^{2},\quad S(x+1,y+1)=S(x,y);$
  2. (2)
    there exists $\varepsilon>0$ such that
    $\frac{\partial^{2}S}{\partial x\partial y}S(x,y)<-\varepsilon,\quad\forall(x,y)\in\mathbb{R}\times\mathbb{R}.$
  3. (3)
    for all $(x,p),(y,p^{\prime})\in\mathbb{R}^{2}$, we have the following equivalence
    $F(x,p)=(y,p^{\prime})\Longleftrightarrow\begin{cases}p=-\partial_{1}S(x,y)\\ p^{\prime}=\partial_{2}S(x,y)\end{cases}.$
Remark 2.18.
For any lift $F$ of a twist map $f$, there exists a unique generating function $S$, up to addition by a constant, that satisfies Definition 2.17 (see [MMS89, Appendix 1]).
One can then state Aubry & Le Daeron Non-crossing Lemma:
Lemma 2.19 ([ALD83]).
Let $f$ be a twist map and $S$ be a generating function. If $x_{1},x_{2},y_{1},y_{2}$ are four reals such that $(x_{1}-x_{2})(y_{1}-y_{2})<0$, then
$S(x_{1},y_{1})+S(x_{2},y_{2})>S(x_{1},y_{2})+S(x_{2},y_{1}).$
This Non-crossing Lemma can be obtained by integrating over a quadrangle and we omit its detailed proof.
Proposition 2.20.
Let $f$ be a twist map. For any $c\in\mathbb{R}$, the function
$S_{c}(x,y):=S(x,y)+c(x-y)$
satisfies Hypothesis 1.
Proof.
Items (4) and (5) of Hypothesis 1 straightforwardly follow from items (1) and (2) of Definition 2.17 respectively. Since $S$ is of class $C^{2}$, we deduce that $S_{c}$ is of class $C^{1}$ and locally semiconcave (see [CS04, Proposition 2.1.2]). Remaining work is to demonstrate that $S_{c}$ is coercive. This conclusion follows from the superlinearity of the generating function $S$, expressed as
$\lim_{|x-y|\rightarrow+\infty}\frac{S(x,y)}{|x-y|}=+\infty.$
It follows immediately from a lemma due to MacKay, Meiss and Stark’s article [MMS89, Lemma A1.6]
Remark 2.21.
  1. (a)
    The function $S_{c}$ is referred to as the generating function at the cohomology class $c$ of $\mathbb{T}$. As proved in Proposition 2.20, all theories in Section 2.1 are applicable to $S_{c}$. For instance, consider any $c$ in the first cohomology group $H^{1}(\mathbb{T},\mathbb{R})$, which is isomorphic to $\mathbb{R}$, the Discrete Weak K.A.M. Theorem 2.7 defines the effective interaction $\bar{S_{c}}$. The function $\alpha(c)=-\bar{S_{c}}$ is called Mather’s $\alpha$-function, which is $C^{1}$ (see [Zav23, Theorem 5.4.9]). As a second example, the Aubry set $\mathcal{A}_{c}^{*}$ at cohomology class $c$ is $F_{c}$-invariant, where
    $F_{c}(x,p)=F(x,p+c)-(0,c).$
    This corresponds to the discrete standard map associated with $S_{c}$. Consequently, the set $(0,c)+\mathcal{A}_{c}^{*}$ is $F$-invariant, where $F$ is the lift of the twist map. Thus, by examining the discrete weak K.A.M. theory of $S_{c}$, we can investigate the dynamics of $F$ at the cohomology class $c$.
  2. (b)
    Since the dimension $d=1$, the the Aubry set $\mathcal{A}_{c}^{*}$ at cohomology class $c$ possesses additional properties beyond Proposition 2.14 (see [GT11, Theorem 11.10]):
    • $(0,c)+\mathcal{A}^{*}_{c}$ is an $F$-ordered set, i.e., $\forall z,z^{\prime}\in(0,c)+\mathcal{A}^{*}_{c}$,
      $\pi_{1}(z)<\pi_{1}(z^{\prime})\quad\Longrightarrow\quad\pi_{1}(F(z))<\pi_{1}(F(z^{\prime})).$
[…]
3. Proof of Theorem 1.1
3.1. $\Sigma_{+}$ is well-defined, non-decreasing and continuous
Let $u\in C^{0}(\mathbb{T})$. Define the set-valued map
$\Sigma_{+}:\mathbb{R}$ $\rightarrow 2^{\mathbb{R}}$
$x$$\mapsto\{y\in\mathbb{R}\mid T^{+}[u](x)=u(y)-S(x,y)\}.$
Thanks to Remark 2.4, we have $\Sigma_{+}(x)$ is non-empty for each $x\in\mathbb{R}$. For two subsets $A\subset\mathbb{R}$ and $B\subset\mathbb{R}$, we define $A\leq B$ if and only if $x_{1}\leq x_{2},\forall x_{1}\in A,x_{2}\in B.$
Lemma 3.1.
If $u\in C^{0}(\mathbb{T})$, then $\Sigma_{+}$ is non-decreasing. That is, if $x_{1}<x_{2}$, then $\Sigma_{+}(x_{1})\leq\Sigma_{+}(x_{2})$.
Proof.
Suppose that $x_{1}<x_{2}$ but $y_{1}>y_{2}$, where $y_{1}\in\Sigma_{+}(x_{1})$ and $y_{2}\in\Sigma_{+}(x_{2})$. We will use Non-crossing Lemma (Lemma 2.19) to reduce a contradiction. We know $(x_{1}-x_{2})(y_{1}-y_{2})<0$ and the Non-crossing Lemma says
$S(x_{1},y_{1})+S(x_{2},y_{2})>S(x_{1},y_{2})+S(x_{2},y_{1}).$
Using the definition of the set-valued map $\Sigma_{+}$, we have that
$u(y_{1})-S(x_{1},y_{1})=T^{+}[u](x_{1})=\sup_{y\in\mathbb{R}}\big{\{}u(y)-S(x_{1},y)\big{\}}\geq u(y_{2})-S(x_{1},y_{2});$
$u(y_{2})-S(x_{2},y_{2})=T^{+}[u](x_{2})=\sup_{y\in\mathbb{R}}\big{\{}u(y)-S(x_{2},y)\big{\}}\geq u(y_{1})-S(x_{2},y_{1}).$
Adding these two inequalities together, we get
$S(x_{1},y_{2})+S(x_{2},y_{1})\geq S(x_{1},y_{1})+S(x_{2},y_{2}).$
This contradicts the Non-crossing Lemma and we complete the proof. ∎
Lemma 3.2.
If $u$ is a discrete weak K.A.M. solution, then
$\bigcup_{x\in\mathbb{R}}\Sigma_{+}(x)=\mathbb{R}.$
Indeed, for any $y\in\mathbb{R}$, there exists $x\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Specifically, $y\in\Sigma_{+}(x)$.
Proof.
Thanks to Remark 2.4, for any $y\in\mathbb{R}$, there exists $x\in\mathbb{R}$ such that
$T_{-}[u](y)=u(x)+S(x,y).$
Since $u$ is a discrete weak K.A.M. solution, we have
$u(y)+\bar{S}=T_{-}[u](y)=u(x)+S(x,y).$
That is, the pair $(x,y)$ is $u$-calibrated. Hence by Remark 2.11, we have $T^{+}[u](x)=u(y)-S(x,y)$ That is, $y\in\Sigma_{+}(x)$. ∎
Proposition 3.3.
If $u$ is a discrete weak K.A.M. solution, then for any $x\in\mathbb{R}$, the set $\Sigma_{+}(x)$ is a singleton.
Proof.
Suppose that there exists $y_{1}<y_{2}$ such that $y_{1},y_{2}\in\Sigma_{+}(x_{0})$. By Lemma 3.1, for any $x>x_{0}$, we have $\Sigma_{+}(x)\geq y_{2}$; for any $x<x_{0}$, we have $\Sigma_{+}(x)\leq y_{1}$. Hence by Lemma 3.2, we have
$(y_{1},y_{2})\subset\Sigma_{+}(x_{0}).$
That is,
$u(y)-S(x_{0},y)=T^{+}[u](x_{0}),\quad\forall y\in(y_{1},y_{2}).$
Since $S$ is $C^{1}$, we have that $u$ is differentiable on $(y_{1},y_{2})$ and
$\ \mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)=\partial_{2}S(x_{0},y),\quad\forall y\in(y_{1},y_{2}).$
Using Theorem 2.15, we have
$(x_{0},-\partial_{1}S(x_{0},y))=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}y)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u),\quad\forall y\in(y_{1},y_{2}).$
Hence
$-\partial_{1}S(x_{0},y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{0}),\quad\forall y\in(y_{1},y_{2}),$
which contradicts $y_{1}\neq y_{2}$ and that $S$ is ferromagnetic. ∎
Using (b) of Lemma 2.6, we have the following corollary.
Corollary 2.
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is differentiable everywhere.
If $u$ is a discrete weak K.A.M. solution, then $\Sigma_{+}:\mathbb{R}\rightarrow\mathbb{R}$ is a well-defined map. It follows from Lemma 3.1 that $\Sigma_{+}$ is non-decreasing as a real-valued map. Since discontinuities of monotonic function are jump discontinuities, but Lemma 3.2 implies that $\Sigma_{+}$ has no jump discontinuity. We conclude that $\Sigma_{+}$ is continuous.
Utilizing the periodicity of $u$ and $T^{+}[u]$, along with the diagonal-periodicity of $S$, we have
$y=\Sigma_{+}(x)$ $\quad\Longleftrightarrow\quad T^{+}[u](x)=u(y)-S(x,y)$
$\quad\Longleftrightarrow\quad T^{+}[u](x+1)=u(y+1)-S(x+1,y+1)$
$\quad\Longleftrightarrow\quad y+1=\Sigma_{+}(x+1).$
This implies $\Sigma_{+}(x+1)=\Sigma_{+}(x)+1$. In other words, $\Sigma_{+}$ is a lift of a circle map of degree $1$.
Remark 3.4.
Notice that Zavidovique’s notes [Zav23, Corollary 5.3.13] states the regularity of $T^{+}[u]$ whenever $u$ is a weak K.A.M. solution.
Proposition 3.5 ([Zav23]).
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is $C^{1}$.
Our method is different from Zavidovique’s because Zavidovique used one theorem due to Arnaud (see [Arn11, Corollary 2]). Furthermore, we will strengthen Zavidovique’s result in Corollary 4 to $T^{+}[u]$ is $C^{1,1}$, provided $u$ is a discrete weak K.A.M. solution.
3.2. $\Sigma_{+}$ is Lipschitz
First we prove a strengthened version of Lemma 3.2.
Proposition 3.6.
Let $u$ be a discrete weak K.A.M. solution. For any $y\in\mathbb{R}$, the preimage $(\Sigma_{+})^{-1}(y)$ is either a singleton or a non-degenerate closed bounded interval. If $(\Sigma_{+})^{-1}(y)$ is a singleton, denoted by $x$, then $(x,y)\in\mathcal{NS}(u)$. If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval, denoted by $[x_{-},x_{+}]$, then $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$.
Proof.
By Lemma 3.2, we know that $(\Sigma_{+})^{-1}(y)$ is nonempty for all $y\in\mathbb{R}$.
Case 1: $(\Sigma_{+})^{-1}(y)$ is a singleton $x$. By Lemma 3.2, we can find $x^{\prime}\in\mathbb{R}$ such that $(x^{\prime},y)\in\mathcal{NS}(u)$ and $x^{\prime}\in(\Sigma_{+})^{-1}(y)$. Hence by the singleton property of $(\Sigma_{+})^{-1}(y)$, we have $x^{\prime}=x$ and $(x,y)\in\mathcal{NS}(u)$.
Case 2: $(\Sigma_{+})^{-1}(y)$ contains at least two different points $x_{1}<x_{2}$. Since $\Sigma_{+}$ is continuous and satisfies $\Sigma_{+}(x+1)=\Sigma_{+}(x)+1$, we claim that
$x_{-}:=\inf\{x\in\mathbb{R}\mid\Sigma_{+}(x)=y\},\quad x_{+}:=\sup\{x\in\mathbb{R}\mid\Sigma_{+}(x)=y\}$
satisfy $-\infty<x_{-}\leq x_{1}<x_{2}\leq x_{+}<+\infty$ and $x_{-},x_{+}\in(\Sigma_{+})^{-1}(y)$. Since $\Sigma_{+}$ is non-decreasing, we have $(\Sigma_{+})^{-1}(y)=[x_{-},x_{+}]$, which is a non-degenerate closed bounded interval.
Before showing $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$, let us first prove that the cardinal of
[…]
is at most countable. We give an injection from $\mathcal{C}$ to $\mathbb{Q}$. We consider the map
$(\Sigma_{+})^{-1}:\mathcal{C}$ $\rightarrow 2^{\mathbb{R}}$
$y$ $\mapsto[x_{-}(y),x_{+}(y)]$
Since $\Sigma_{+}$ is a well-defined map, for $y_{1}\neq y_{2}$, we have
$[x_{-}(y_{1}),x_{+}(y_{1})]\cap[x_{-}(y_{2}),x_{+}(y_{2})]=\emptyset.$
By Axiom of Choice, we can pick a rational number $q(y)$ from each $[x_{-}(y),x_{+}(y)]$. Hence for $y_{1}\neq y_{2}$, we get $q(y_{1})\neq q(y_{2})$. This defines an injection from $\mathcal{C}$ to $\mathbb{Q}$, whose cardinal is countable.
If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval $[x_{-},x_{+}]$, then we can find an increasing sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $x_{n}\rightarrow x_{-}$ and
$(\Sigma_{+})^{-1}(y_{n}),\quad\text{ where }y_{n}=\Sigma_{+}(x_{n}),$
is a singleton for all $n\in\mathbb{Z}_{+}$. Suppose this is incorrect, then there exists $\delta>0$ such that for all $x\in[x_{-}-\delta,x_{-}]$, $(\Sigma_{+})^{-1}(\Sigma_{+}(x))$ is not a singleton (so it must be a non-degenerate closed interval). Since $x_{-}=\inf\{x\mid\Sigma_{+}(x)=y\}$, we have $\Sigma_{+}(x_{-}-\delta)<y$. Now using the fact that $\mathcal{C}$ is at most countable, we claim there exists $y^{\prime}$ such that $\Sigma_{+}(x_{-}-\delta)<y^{\prime}<y$ and $(\Sigma_{+})^{-1}(y^{\prime})$ is a singleton, which is denoted by $x^{\prime}$. Since $\Sigma_{+}$ is non-decreasing, we have $x^{\prime}\in[x_{-}-\delta,x_{-}]$, but $(\Sigma_{+})^{-1}(y^{\prime})$ is a singleton. Contradiction!
Thus by Case 1, the sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ satisfies $(x_{n},y_{n})\in\mathcal{NS}(u)$. Since $\mathcal{NS}(u)$ is closed and $\lim_{n\to+\infty}(x_{n},y_{n})=(x_{-},y)$, we have $(x_{-},y)\in\mathcal{NS}(u)$. By a similar argument, we similarly obtain $(x_{+},y)\in\mathcal{NS}(u)$. ∎
Corollary 3.
Let $u$ be a discrete weak K.A.M. solution. For any $x\in\mathbb{R}$, there exist $x^{\prime},x^{\prime\prime}\in\mathbb{R}$ such that $x^{\prime}\leq x\leq x^{\prime\prime}$, $\Sigma_{+}(x^{\prime})=\Sigma_{+}(x)=\Sigma_{+}(x^{\prime\prime})$ and $x^{\prime},x^{\prime\prime}\in\pi_{1}(\mathcal{NS}(u))$.
Proof.
Take $x\in\mathbb{R}$ and denote $\Sigma_{+}(x)$ by $y$. If $(\Sigma_{+})^{-1}(y)$ is a singleton $x$, then take $x^{\prime}=x^{\prime\prime}=x$. If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed interval $[x_{-},x_{+}]$, then take $x^{\prime}=x_{-}$ and $x^{\prime\prime}=x_{+}$. ∎
Lemma 3.7.
Let $u$ be a discrete weak K.A.M. solution. If $x\in\pi_{1}(\mathcal{NS}(u))$, then
  1. (i)
    $u(x)-\bar{S}=T^{+}[u](x)$;
  2. (ii)
    Both $u$ and $T^{+}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)$.
Proof.
Let $x\in\pi_{1}(\mathcal{NS}(u))$, then there exists $y\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Hence $u(x)-\bar{S}=T^{+}[u](x)$ by Remark 2.11. Pick $w\in(\Sigma_{+})^{-1}(x)$, i.e., $T^{+}[u](w)=u(x)-S(w,x)$. Combine these two equations and we get
$T^{+}[u](x)-T^{+}[u](w)=\left(u(x)-\bar{S}\right)-\left(u(x)-S(w,x)\right)=S(w,x)-\bar{S}.$
That is, $(w,x)\in\mathcal{NS}(T^{+}[u])$. Hence by (b) of Lemma 2.12, we have both $u$ and $T^{+}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)$. ∎
Lemma 3.8.
If $u$ is a discrete weak K.A.M. solution, then $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u$ is Lipschitz on $\pi_{1}(\mathcal{NS}(u))$. That is, there exists $K>0$ such that
$|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})|\leq K|x_{1}-x_{2}|,\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Proof.
Since $u$ is a semiconcave function, by (i) of Proposition 2.2, there exists a constant $C\geq 0$ such that
$u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. On the other hand, since $T^{+}[u]$ is semiconvex, there exists $C^{\prime}\geq 0$ such that
$T^{+}[u](x_{1})-T^{+}[u](x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{2})(x_{1}-x_{2})\geq-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. Using Lemma 3.7 to replace $T^{+}[u]$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ with $u$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u$ respectively, we have
$u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\geq-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. Therefore we have
$-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}\leq u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Exchange the roles of $x_{1}$ and $x_{2}$, then we get
$-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}\leq u(x_{2})-u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})(x_{2}-x_{1})\leq\frac{C}{2}|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Add these two inequalities together, we get
$-C^{\prime}|x_{1}-x_{2}|^{2}\leq(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))(x_{1}-x_{2})\leq C|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Hence if $K=\max\{C,C^{\prime}\}$, then
$|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})|\leq K|x_{1}-x_{2}|,\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).\qquad\qquad\qed$
Proposition 3.9.
If $u$ is a discrete weak K.A.M. solution, then $\Sigma_{+}$ is Lipschitz.
Proof.
For any $x\in\pi_{1}(\mathcal{NS}(u))$, there exists $y\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Since $u$ is a weak K.A.M. solution, we have $(x,y)\in\mathcal{NS}(T^{-}[u])$. By (a) of Lemma 2.12, we know $u$ is differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=-\partial_{1}S(x,y)$. Since $S$ is ferromagnetic, we have $y=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))$, where $F$ is the lift of $f$ associated to $S$. On the other hand, By $(x,y)\in\mathcal{NS}(u)$ and Remark 2.11, we have $y=\Sigma_{+}(x)$. Therefore, we have
$\Sigma_{+}(x)=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)),\quad\forall x\in\pi_{1}(\mathcal{NS}(u)).$
Then we claim that $\Sigma_{+}$ is also Lipschitz on $\pi_{1}(\mathcal{NS}(u))$. In fact, for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$,
$|\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2})|$ $=\left|\pi_{1}\circ F(x_{1},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1}))-\pi_{1}\circ F(x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))\right|$
$\leq K^{\prime}\|(x_{1}-x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))\|_{\mathbb{R}^{2}}$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}-x_{2}|.$
where
$K^{\prime}=\sup_{x\in[0,1],\ |p|\leq K}\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}(\pi_{1}\circ F)(x,p)\|_{(\mathbb{R}^{2})^{*}}.$
Now we show that $\Sigma_{+}$ is Lipschitz on whole $\mathbb{R}$. For any $x_{1},x_{2}\in\mathbb{R}$, by Corollary 3, there exist $x_{1}^{\prime\prime},x_{2}^{\prime}\in\pi_{1}(\mathcal{NS}(u))$ such that
$x_{1}\leq x_{1}^{\prime\prime}\leq x_{2}^{\prime}\leq x_{2},\quad\Sigma_{+}(x_{1}^{\prime\prime})=\Sigma_{+}(x_{1})\quad\text{ and }\quad\Sigma_{+}(x_{2}^{\prime})=\Sigma_{+}(x_{2}).$
Then
$|\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2})|$ $=|\Sigma_{+}(x_{1}^{\prime\prime})-\Sigma_{+}(x_{2}^{\prime})|$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}^{\prime\prime}-x_{2}^{\prime}|$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}-x_{2}|.$
Hence $\Sigma_{+}$ is Lipschitz. ∎
Proposition 3.10.
Let $u$ be a weak K.A.M. solution. If $F$ is the lift of $f$ associated to $S$, then
$\forall x\in\mathbb{R},\quad\Sigma_{+}(x)=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
Proof.
Take $x\in\mathbb{R}$ and let $y=\Sigma_{+}(x)$. Then $T^{+}[u](x)=u(y)-S(x,y)$. Using (b) of Lemma 2.5, we have that
$-\partial_{1}S(x,y)\in\nabla^{-}T^{+}[u](x).$
Since $T^{+}[u]$ is differentiable, by (iv) if Proposition 2.2, we know that $\nabla^{-}T^{+}[u](x)=\{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)\}.$ Hence we get
$-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x).$
Since $S$ is ferromagnetic, if $F$ is the lift of $f$ associated to $S$, then
$y=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
The proof ends. ∎
Corollary 4.
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is a $C^{1,1}$ function.
Proof.
We only need to show that $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ is Lipschitz. Using Proposition 3.10 and the ferromagnetic condition of $S$, we have
$\ \mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)=-\partial_{1}S(x,\Sigma_{+}(x)).$
Then for any $x_{1},x_{2}\in\mathbb{R}$,
$\left|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{2})\right|$ $=\left|\partial_{1}S(x_{1},\Sigma_{+}(x_{1}))-\partial_{1}S(x_{2},\Sigma_{+}(x_{2}))\right|$
$\leq K^{\prime\prime}\|(x_{1}-x_{2},\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2}))\|_{\mathbb{R}^{2}}$
$\leq K^{\prime\prime}\sqrt{1+K^{\prime 2}+(K^{\prime}K)^{2}}|x_{1}-x_{2}|,$
where
$K^{\prime\prime}=\sup_{x\in[0,1],\ |y|\leq\|\Sigma_{+}\|_{C^{0}[0,1]}}\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}(\partial_{1}S)(x,y)\|_{(\mathbb{R}^{2})^{*}}.$
Hence $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ is Lipschitz. ∎
3.3. $\Sigma_{+}$ propagates singularities
Proposition 3.11.
Let $u$ be a weak K.A.M. solution. Then $y\in\operatorname{Sing}(u)$ if and only if $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval.
Proof.
Since $u$ is a weak K.A.M. solution, we know $u=T^{-}u-\bar{S}$. If $y\in\operatorname{Sing}(u)$, then $T^{-}[u]$ is also non-differentiable at $y$. Using (a) of Lemma 2.6, we have that there exist $x_{1}\neq x_{2}$ such that $T^{-}[u](y)=u(x_{i})+S(x_{i},y),\ \forall i\in\{1,2\}$. Using that $u$ is a weak K.A.M. solution again, we claim
$u(y)+\bar{S}=T^{-}[u](y)=u(x_{i})+S(x_{i},y),\ \forall i\in\{1,2\}.$
That is, $(x_{i},y)\in\mathcal{NS}(u),\ \forall i\in\{1,2\}.$ Hence $T^{+}[u](x_{i})=u(y)-S(x_{i},y)$ by Remark 2.11. That is, $\{x_{1},x_{2}\}\subset(\Sigma_{+})^{-1}(y)$. By Proposition 3.6, $(\Sigma)^{-1}(y)$ must be a non-degenerate closed bounded interval.
If $(\Sigma)^{-1}(y)$ is a non-degenerate closed bounded interval $[x_{-},x_{+}]$, then by Proposition 3.6, we know that $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$. Hence $T^{-}[u](y)=u(x_{\pm})+S(x_{\pm},y)$ by Remark 2.11. By (a) of Lemma 2.6, we have $T-[u]$ is non-differentiable at $y$. Hence $y\in\operatorname{Sing}(u)$. ∎
Proposition 3.12.
Let $u$ be a weak K.A.M. solution. If $x\in\operatorname{Sing}(u)$, then $\Sigma_{+}(x)\in\operatorname{Sing}(u)$.
Proof.
Let $y=\Sigma_{+}(x)$. By Corollary 3, there exists $x^{\prime}$ such that $\Sigma_{+}(x^{\prime})=y$ and $(x^{\prime},y)\in\mathcal{NS}(u)$. Since $u$ is a weak K.A.M. solution, we know $u=T^{-}u-\bar{S}$. Hence $(x^{\prime},y)\in\mathcal{NS}(T^{-}[u])$ and it follows from (a) of Lemma 2.12 that $u$ is differentiable at $x^{\prime}$. Then $x\neq x^{\prime}$ since $x\in\operatorname{Sing}(u)$. Hence we claim that $(\Sigma)^{-1}(y)$ is not a singleton because it contains at least $x^{\prime}$ and $x$ By Proposition 3.6, $(\Sigma)^{-1}(y)$ must be a non-degenerate closed bounded interval. By Proposition 3.11, we have $y\in\operatorname{Sing}(u)$. ∎
3.4. $\Sigma_{+}$ admits a rotation number
The following lemma is classical. The reader may refer to any textbook on rotation theory. For the sake of self-containedness, we provide the proof here. We do not make any claim of originality for this lemma.
Lemma 3.13.
Let Φ:normal-:normal-$\Phi:\mathbb{R}\rightarrow\mathbb{R}$ be a continuous map such that
  1. (i)
    $\Phi(x+1)=\Phi(x)+1$, for any $x\in\mathbb{R}$;
  2. (ii)
    if $x_{1}<x_{2}$, then $\Phi(x_{1})\leq\Phi(x_{2})$.
Then the limit
$\rho(\Phi):=\lim_{n\to+\infty}\frac{1}{n}\left(\Phi^{n}(x)-x\right)$
exists for all $x\in\mathbb{R}$ and is independent of $x$.
Proof.
Using the property (i), by induction, one can show
$\forall x\in\mathbb{R},m\in\mathbb{N}_{+},\quad\Phi^{m}(x+1)=\Phi^{m}(x)+1.$
Hence $\Phi^{m}-\operatorname{id}$ is $1$-periodic and by continuity of $\Phi$ we can denote
$\alpha_{m}:=\inf_{\mathbb{R}}\{\Phi^{m}-\operatorname{id}\},\quad\quad\beta_{m}:=\sup_{\mathbb{R}}\{\Phi^{m}-\operatorname{id}\}.$
Using the property (ii), we claim that
$0\leq\beta_{m}-\alpha_{m}\leq 1$
In fact, for any $x_{1}\leq x_{2}\leq x_{1}+1$, we have that $\Phi^{m}(x_{1})\leq\Phi^{m}(x_{2})\leq\Phi^{m}(x_{1})+1$. Then
$\Phi^{m}(x_{1})-(x_{1}+1)\leq\Phi^{m}(x_{2})-x_{2}\leq\Phi^{m}(x_{1})+1-x_{1}.$
Hence $|(\Phi^{m}(x_{1})-x_{1})-(\Phi^{m}(x_{2})-x_{2})|\leq 1$ and thus $\beta_{m}-\alpha_{m}\leq 1$.
For any $n\in\mathbb{N}_{+},n\geq m$, we know that
$n=qm+r,\quad 0\leq r<m.$
It follows from $\forall y\in\mathbb{R},\ \alpha_{m}\leq\Phi^{m}(y)-y\leq\beta_{m},$ that
$\forall p\in\mathbb{N}_{+},\quad\alpha_{m}\leq\Phi^{pm}(y)-\Phi^{(p-1)m}(y)\leq\beta_{m}.$
Adding these inequalities for $p=1,2,\dots,q$, we get
$q\alpha_{m}\leq\Phi^{qm}(y)-y\leq q\beta_{m}.$
In particular, we have that
$\begin{cases}q\alpha_{m}\leq\Phi^{qm+r}(x)-\Phi^{r}(x)\leq q\beta_{m};\\ r\alpha_{1}\leq\Phi^{r}(x)-x\leq r\beta_{1}.\end{cases}$
Adding these two inequalities together and dividing them by $n$, we get
$\frac{q\alpha_{m}+r\alpha_{1}}{n}\leq\frac{1}{n}\left(\Phi^{n}(x)-x\right)\leq\frac{q\beta_{m}+r\beta_{1}}{n}.$
Since
$\left|\frac{q\beta_{m}+r\beta_{1}}{n}-\frac{q\alpha_{m}+r\alpha_{1}}{n}\right|$ $\rightarrow\frac{\beta_{m}-\alpha_{m}}{m}\quad(n\to+\infty),$
$\frac{\beta_{m}-\alpha_{m}}{m}$ $\rightarrow 0\quad\quad\quad\quad\ \ (m\to+\infty),$
[…]
By combining Section 3.1 with Lemma 3.13, we proved that $\Sigma_{+}$ admits a rotation number $\rho(\Sigma_{+})$ for positive iterations. Moreover, for any $(x,p)\in\mathcal{A}^{*}$, by Proposition 3.10 we have
$\Sigma_{+}(x)=\pi_{1}\circ F(x,p).$
Hence $\rho(\Sigma_{+})$ equals to the rotation number of the Aubry set.
4. An explanation of Arnaud’s observation
In this section, we aim to prove Theorem 1.3. We first provide a formula of $(\Sigma_{+})^{-1}$.
Proposition 4.1.
Let $u$ be a discrete weak K.A.M. solution. If $F$ is the lift of $f$ associated to $S$, then
$\forall y\in\mathbb{R},\quad(\Sigma_{+})^{-1}(y)=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).$
Proof.
Case 1: $u$ is differentiable at $y$. By Proposition 3.11, we know that $(\Sigma_{+})^{-1}(y)=\{x\}$ is a singleton. Using (b) of Lemma 2.5, we claim $\partial_{2}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)$. Hence the definition of generating function implies $x=\pi_{1}\circ F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
Case 2: $u$ is non-differentiable at $y$. By Proposition 3.11, we know that $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed interval $[x_{-},x_{+}]$. According to the proof of Proposition 3.6, we can find an increasing sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $x_{n}\rightarrow x_{-}$ and $u$ is differentiable at $y_{n}:=\Sigma_{+}(x_{n})$ for all $n\in\mathbb{Z}_{+}$. Then
$x_{n}=\pi_{1}\circ F^{-1}(y_{n},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y_{n}))$
by Case 1 and we claim that
$x_{-}\in\pi_{1}\circ F^{-1}(y,\nabla^{*}u(y)).$
Similarly, we also have $x_{+}\in\pi_{1}\circ F^{-1}(y,\nabla^{*}u(y))$. By (v) of Proposition 2.2 and item (a) of Remark 2.4, *$\nabla^{*}u(y)\subset\partial\nabla^{+}u(y)$ and then we have that
$x_{-},x_{+}\in\pi_{1}\circ F^{-1}(y,\partial\nabla^{+}u(y)).$
Since $\nabla^{+}u(y)$ is a closed convex subset of $\mathbb{R}$, its boundary contains at most two different points, that is,
$\{x_{-},x_{+}\}=\pi_{1}\circ F^{-1}(y,\partial\nabla^{+}u(y)).$
Then noticing $x_{-}<x_{+}$ and using the twist condition of $F$, we claim that
$[x_{-},x_{+}]=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).$
The proof is complete. ∎
Now combining Propositions 3.10 and 4.1, we can manually verify the bijectivity of the commutative diagram below:
{tikzcd}
where $\Gamma^{+}(x,y)=(y,\partial_{2}S(x,y))$, $\Gamma^{-}(x,y)=(x,-\partial_{1}S(x,y))$ and $F$ is the lift of $f$ associated to $S$. We only show that the map $\Gamma^{-}$ in the above diagram is bijective, and the bijectivity of $\Gamma^{+}$ and $F$ could be proved similarly. That is, for any $x\in\mathbb{R}$,
$\Gamma^{-}(x,\Sigma_{+}(x))$ $=(x,-\partial_{1}S(x,\Sigma_{+}(x)))$
$=(x,-\partial_{1}S(x,\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x))))$
$=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
Because the twist condition, which says all $\Gamma^{+},\Gamma^{-}$ and $F$ are $C^{1}$ diffeomorphisms, and the notion of Lipschitz submanifolds is invariant by $C^{1}$ diffeomorphisms, it follows from Proposition 3.9 that all these three graphs $\operatorname{Graph}(\Sigma_{+}),\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])$ and $\operatorname{Graph}(\nabla^{+}u)$ are Lipschitz submanifolds of $\mathbb{R}^{2}$. This finishes the proof of Theorem 1.3.
5. Singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$
In this section, our goal is to analyze the singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$ and to provide proofs of Theorems 1.4 and 1.5. Indeed, Theorem 1.4 follows directly from Theorem 1.1. It suffices to show Theorem 1.5.
Lemma 5.1.
Let $u$ be a discrete weak K.A.M. solution. If $u$ is differentiable at $y$, then
$(\Lambda_{0})^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
Proof.
Denote $(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)):=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))$. Using Lemma 3.2, there exists $x^{\prime}$ such that $(x^{\prime},y)\in\mathcal{NS}(u)$. Then it follows from Remark 2.11 and item (a) of Lemma 2.5 that $\partial_{2}S(x^{\prime},y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)$. Since $S$ is ferromagnetic, we have $x=x^{\prime}$ and $(x,y)\in\mathcal{NS}(u)$. Hence $\Sigma_{+}(x)=y$. It follows from $y\not\in\operatorname{Sing}(u)$ and Proposition 3.11 that $x=(\Sigma_{+})^{-1}(y)$. Hence $(\Lambda_{0})^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))$. ∎
Notice that once we prove item (ii) of Theorem 1.5, then the item (i) of Theorem 1.5 can be obtained directly from Theorem 2.15.
Proof of (ii) in Theorem 1.5.
Since *$\mathcal{I}^{*}(u)$ is $F$ invariant, for any $(x,p)\in\mathcal{I}^{*}(u)$, we have $(y,p^{\prime}):=F(x,p)\in\mathcal{I}^{*}(u)$. Then $u$ is differentiable at $y$. Using Lemma 5.1, we have
$(x,p)=F^{-1}(y,p^{\prime})=(\Lambda_{0})^{-1}(y,p^{\prime}).$
Hence, $\Lambda_{0}(x,p)=(y,p^{\prime})=F(x,p)$. That is, the restriction of $\Lambda_{0}$ on *$\mathcal{I}^{*}(u)$ equals $F$. ∎
The item (iii) of Theorem 1.5 follows immediately from item (iv) of Theorem 1.4 and the fact $\operatorname{Sing}(u)$ is at most countable (see [CS04, Proposition 4.1.3]).
Proof of (iv) in Theorem 1.5.
Fix $(x,p)\in\operatorname{Graph}(\nabla^{+}u)$. Suppose that for any $n\in\mathbb{Z}_{+}$, we have
$\Lambda_{0}^{n}(x,p)\not\in\mathcal{K}^{*}(u).$
Then by the definition of $\mathcal{K}^{*}(u)$, for any $n\in\mathbb{Z}_{+}$, we have
$\pi_{1}\circ\Lambda_{0}^{n}(x,p)\not\in\operatorname{Sing}(u).$
Using Lemma 5.1, by induction, we have
$\Lambda_{0}^{n}(x,p)=F^{n}(x,p),\quad\forall n\in\mathbb{Z}_{+}.$
Therefore, for any $n\in\mathbb{Z}_{+}$, we have
$F^{n}(x,p)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
That is, $(x,p)\in\mathcal{I}^{*}(u)$. ∎
Appendix
For selfcontainedness, we give proofs for Lemmas 2.5, 2.6 and 2.12.
Proof of Lemma 2.5.
(a). Suppose that $T^{-}[u](y)=u(x)+S(x,y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$.
To establish the first inclusion, $-\partial_{1}S(x,y)\in\nabla^{-}u(x)$, we begin with the definition of the discrete backward Lax-Oleinik operator : $T^{-}[u](y)=\inf_{z\in\mathbb{R}^{d}}u(z)+S(z,y)$. Then we have
$u(x)+S(x,y)\leq u(z)+S(z,y),\ \quad\forall z\in\mathbb{R}^{d}.$
Let $\Omega$ be a bounded convex open neighborhood of $x$. Since $S$ is locally semiconcave, we have $S(\cdot,y)$ is semiconcave on $\Omega$. By utilizing items (i) and (iv) of Proposition 2.2, there exists $C\geq 0$ such that
$S(z,y)-S(x,y)\leq\partial_{1}S(x,y)(z-x)+\frac{C}{2}\|z-x\|^{2}$
for any $z\in\Omega$. Combining these two inequalities, we get
$u(x)-u(z)\leq\partial_{1}S(x,y)(z-x)+\frac{C}{2}\|z-x\|^{2},\ \quad\forall z\in\Omega.$
Hence, $\partial_{1}S(x,y)\in\nabla^{+}(-u)(x)$ follows from item (ii) of Proposition 2.2. Since $\nabla^{+}(-u)(x)=-\nabla^{-}u(x)$, we conclude that $-\partial_{1}S(x,y)\in\nabla^{-}u(x)$.
Now, we move on to the second inclusion, $\partial_{2}S(x,y)\in\nabla^{+}T^{-}[u](y)$. According to the definition of the discrete backward Lax-Oleinik operator $T^{-}$, we have $T^{-}[u](\cdot)\leq u(x)+S(x,\cdot)$. Hence it follows from items (iii) and (iv) of Proposition 2.2 that
$\{\partial_{2}S(x,y)\}=\nabla^{+}(u(x)+S(x,\cdot))(y)\subset\nabla^{+}T^{-}[u](y).$
(b). The proof for item (b) follows from a similar argument as that of (a). ∎
Proof of Lemma 2.6.
(a). Suppose there exists $x_{1}\neq x_{2}$ such that $T^{-}[u](y)=u(x_{1})+S(x_{1},y)$ and $T^{-}[u](y)=u(x_{2})+S(x_{2},y)$. Using (a) of Lemma 2.5, we get:
$\{\partial_{2}S(x_{1},y),\partial_{2}S(x_{2},y)\}\subset\nabla^{+}T^{-}[u](y).$
The ferromagnetic condition of $S$ asserts that if $x_{1}\neq x_{2}$, then $\partial_{2}S(x_{1},y)\neq\partial_{2}S(x_{2},y)$. Hence $\nabla^{+}T^{-}[u](y)$ is not a singleton set. It follows from item (iv) of Proposition 2.2 that $T^{-}[u]$ is not differentiable at $y$.
Thanks to item (c) of Remark 2.4, we only need to demonstrate that if $x$ is the unique point such that $T^{-}[u](y)=u(x)+S(x,y)$, then $T^{-}[u]$ is differentiable at $y$.
For any converging sequence $y_{k}\rightarrow y$, let us pick $x_{k}$ such that $T^{-}[u](y_{k})=u(x_{k})+S(x_{k},y_{k})$. Since $x$ is the unique point such that $T^{-}[u](y)=u(x)+S(x,y)$, we have $x_{k}\rightarrow x$ due to the continuity of functions $T^{-}[u]$, $u$ and $S$. Then
$T^{-}[u](y_{k})$ $=u(x_{k})+S(x_{k},y)+\left(S(x_{k},y_{k})-S(x_{k},y)\right)$
$\geq T^{-}[u](y)+\left(S(x_{k},y_{k})-S(x_{k},y)\right)$
$=T^{-}[u](y)+\partial_{2}S(x_{k},y)(y_{k}-y)+o(\|y_{k}-y\|)$
Hence $\partial_{2}S(x,y)\in\nabla^{-}T^{-}[u](y)$ by the definition of superdifferential. Using item (a) of Lemma 2.5 and item (iv) of Proposition 2.2, we have that $T^{-}[u]$ is differentiable at $y$.
(b). The proof for item (b) follows from a similar argument as that of (a). ∎
Proof of Lemma 2.12.
(a). If $(x,y)\in\mathcal{NS}(T^{-}[u])$, then by the definition, we have
$T^{-}[u](y)-T^{-}[u](x)=S(x,y)-\bar{S}.$
That is,
$\left[T^{-}[u](y)-u(x)-S(x,y)\right]+\left[u(x)-T^{-}[u](x)+\bar{S}\right]=0.$
From the definitions of $T^{-}$ and sub-actions, we see that
$T^{-}[u](y)-u(x)-S(x,y)\leq 0,\ \quad\ u(x)-T^{-}[u](x)+\bar{S}\leq 0$
respectively. Therefore,
$T^{-}[u](y)-u(x)-S(x,y)=0,\ \quad\ u(x)-T^{-}[u](x)+\bar{S}=0.$
Using (a) of Lemma 2.5, since $T^{-}[u](y)-u(x)-S(x,y)=0$, we have that
$-\partial_{1}S(x,y)\in\nabla^{-}u(x).$
By (c) of Remark 2.4, one can find $w\in\mathbb{R}^{d}$ such that $T^{-}[u](x)=u(w)+S(w,x).$ Using (a) of Lemma 2.5, we know that
$\partial_{2}S(w,x)\in\nabla^{+}T^{-}[u](x).$
Since $u$ is a sub-action, i.e., $u+\bar{S}\leq T^{-}[u]$, and $u(x)+\bar{S}=T^{-}[u](x)$, it follows from the item (iii) of Proposition 2.2 that
$\nabla^{+}u(x)\supset\nabla^{+}T^{-}[u](x)\quad\text{and}\quad\nabla^{-}u(x)\subset\nabla^{-}T^{-}[u](x).$
In conclusion, we have
$-\partial_{1}S(x,y)\in\nabla^{-}u(x)\subset\nabla^{-}T^{-}[u](x);$
$\partial_{2}S(w,x)\in\nabla^{+}T^{-}[u](x)\subset\nabla^{+}u(x).$
By (iv) of Proposition 2.2, we claim that both $u$ and $T^{-}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=dT^{-}[u](x)=-\partial_{1}S(x,y)$.
(b). The proof for item (b) follows from a similar argument as that of (a). ∎